How many scores are above 116?



1
Expert's answer
2022-03-01T08:12:10-0500

Given a population of 5000 scores with a mean of 86 and a standard deviation of 10, how many scores are above 116?


"P(X>116)=1-P(X\le 116)"

"=1-P(Z\le \dfrac{116-86}{10})=1-P(Z\le3)"

"=0.00135"

"0.00135(5000)=7"

7 scores are above 116.